# 2]organization=Shanghai Blockchain Engineering Research Center, city=Shanghai, postcode=200433, country=China

### v1

@@ Abstract
Vertical Federated Learning (VFL) facilitates collaborative machine learning without the need for participants to share raw private data. However, recent studies have revealed privacy risks where adversaries might reconstruct sensitive features through data leakage during the learning process. Although data reconstruction methods based on gradient or model information are somewhat effective, they reveal limitations in VFL application scenarios. This is because these traditional methods heavily rely on specific model structures and/or have strict limitations on application scenarios. To address this, our study introduces the Unified InverNet Framework into VFL, which yields a novel and flexible approach (dubbed UIFV) that leverages intermediate feature data to reconstruct original data, instead of relying on gradients or model details. The intermediate feature data is the feature exchanged by different participants during the inference phase of VFL. Experiments on four datasets demonstrate that our methods significantly outperform state-of-the-art techniques in attack precision. Our work exposes severe privacy vulnerabilities within VFL systems that pose real threats to practical VFL applications and thus confirms the necessity of further enhancing privacy protection in the VFL architecture.
@@ 1 Introduction
In the current field of artificial intelligence, Vertical Federated Learning (VFL) is regarded as a significant transformative strategy Khan et al. (2022); Chen et al. (2020); Tan et al. (2022). VFL allows different institutions to collaborate on model training while protecting their respective data privacy. This approach is particularly applicable in scenarios where different institutions hold different feature data of the same set of users, such as in the finance and healthcare sectors Li et al. (2019); Lu et al. (2020). The participants in a VFL framework consist of an active party and some passive parties Kang et al. (2022). Each passive party owns a set of data features that are fed into a model (called a bottom model) for local training. The active party holds the label information and a top model in addition to its own feature set and bottom model. The active party coordinates the training process by concatenating the bottom model outputs as the input to the top model. This structure enables collaborative model training without the need to share original data, thus preserving the privacy of sensitive data on the participants.
Despite the advantages of VFL in protecting private data, recent studies have shown that it may still face risks of data privacy breaches Geiping et al. (2020); Wu et al. (2022), especially through data reconstruction attacks. Such attacks reconstruct the original features of the training dataset by analyzing intermediate data during the VFL process, potentially leading to sensitive information leaks.
Figure 1 shows a VFL application scenario that is vulnerable to data reconstruction attacks Chen et al. (2022). In this scenario, a bank and a fintech company participate in a VFL for credit analysis, with each entity possessing a subset of user attributes. An assessment agency uses the labels of users to coordinate the training of the VFL model. However, the assessment agency wants to acquire the fintech company’s private data and take it for its own use, so it launches a data reconstruction attack on the model owned by the fintech company. Without violating the VFL protocol, it uses the intermediate data from the VFL to reconstruct the private data owned by the bank, leading to the leakage of the fintech company’s customer information and posing a significant security risk to the actual application of VFL.
@@
Existing data reconstruction attacks on VFL, such as the generative regression network (GRN) method Luo et al. (2021) and the gradient-based inversion attack (GIA) method Jiang et al. (2022), largely draw from methods used in HFL (Horizontal Federated Learning), with a core focus on utilizing model information. However, these methods have certain limitations. For instance, the GRN method is designed for a white-box attack scenario where the reconstruction process requires access to the passive party’s model, which significantly restricts its applicability. Although the GIA method enables a black-box attack, it imposes very strict requirements on the model, demanding a logistic regression model and forbids non-linear activation functions in the output layer of neural networks, which also greatly limits the method’s applicability.
To gain a more thorough understanding of the potential threats to data privacy in VFL, it is crucial to go beyond the existing attack methods and explore a broader range of attack scenarios. The goal of this paper is to fully consider real-world attack scenarios in practical VFL environments and develop effective attack methods for different scenarios. To accomplish this, we have completely abandoned the traditional approach that relies on gradient information or model information and instead have opted to directly utilize the intermediate feature data generated in the VFL framework for attacks. This method constructs an inverse net (InverNet) to effectively extract original data information from the intermediate features output from the target’s model. Following this strategy, We have developed an innovative attack framework, called Unified InverNet Framework in VFL (UIFV), that is applicable to a variety of VFL scenarios with different adversary capabilities. UIFV overcomes the dependency on gradient information or model information of traditional attacks in VFL, thereby providing a more flexible and effective means for data reconstruction attacks in complex VFL environments.
@@
  • Innovative Attack Method: We propose a new type of VFL data reconstruction attack method that does not rely on gradient information or model information but leverages the intermediate features output from the target’s model through an InverNet to reconstruct original data.
@@ 2.1 Vertical Federated Learning
Within the framework of federated learning, based on the distribution characteristics of local data, it is mainly divided into two scenarios: Horizontal Federated Learning (HFL) and Vertical Federated Learning (VFL). In HFL Li et al. (2021b); Zhuang et al. (2021); Li et al. (2021a), the local datasets of data owners have almost no intersection in the sample space, but there is a significant overlap in the feature space. In contrast, in VFL Khan et al. (2022); Chen et al. (2020); Tan et al. (2022), local datasets have a large intersection in the sample space, but little overlap in the feature space. Each of these federated learning types has its applicable scenarios and advantages. VFL is particularly suitable for situations where different institutions hold different feature data of the same set of users; for example, in the financial sector, one institution may have the credit history of users, while another may have their transaction data. Through VFL, these institutions can collaborate to build more accurate risk assessment models without the need to directly exchange sensitive data. VFL has found wide applications in fields such as finance and healthcare Li et al. (2022); Wei et al. (2023).
The participants in a VFL framework consist of an active party and some passive parties Kang et al. (2022). Each passive party owns a set of data features that are fed into a model (called a bottom model) for local training. The active party holds the label information and a top model in addition to its own feature set and bottom model. The active party coordinates the training process by concatenating the bottom model outputs as the input to the top model. This structure enables collaborative model training without the need to share original data, thus preserving the privacy of sensitive data on the participants.
@@ 2.2 Data Reconstruction Attack
Gradient-based methods use gradients generated during machine learning model training for data reconstruction. Beginning with the seminal work Zhu et al. (2019), a substantial body of research, including Geiping et al. (2020) and Zhao et al. (2020), has utilized gradient data for reconstruction, achieving significant empirical results. Vero et al. (2022) introduced TabLeak, a novel method for tabular data, while Pan et al. (2022) investigated the security boundaries of gradient data reconstruction. However, these methods, mainly applicable to HFL architecture, face challenges in VFL architecture where accessing gradient information is difficult. Jin et al. (2021) considered a VFL scenario with shared parameters among underlying models, enabling direct data reconstruction via gradients, but this is not universally applicable in VFL.
Model information-based reconstruction methods use the internal parameters of machine learning models to reconstruct original training data. Luo et al. (2021) proposed the GRN method for recovering data of the passive party in VFL, requiring access to the passive party’s model and its parameters. Jiang et al. (2022) improved the GRN method and proposed GIA, which constructs a shadow model using a small amount of known auxiliary data and their confidence scores, emulating the real passive party’s model for data reconstruction without direct access. Essentially, this approach still relies on utilizing the model’s internal parameters, employing a white-box method for data reconstruction. However, obtaining the passive party’s model in VFL architecture is difficult due to participants’ concerns over data privacy and security, often making them reluctant to share their model details.
Feature-based methods reconstruct data using outputs like Shapley values or intermediate features. In traditional machine learning, Luo et al. (2022) first revealed the risks of feature inference attacks in model explanations based on Shapley values, demonstrating that current explanation methods are prone to privacy leaks. In split learning, He et al. (2019) and He et al. (2020) studied data privacy leakage, and Yin et al. (2023) introduced the Ginver attack, confirming the severity of privacy issues in such systems. Pasquini et al. (2021) explored the risks of malicious servers in learning processes. Despite numerous studies in this area, specific methods for VFL architecture are lacking. Our method reconstructs data using intermediate features of the VFL passive party’s model, treating it as a black-box, significantly broadening the application scenarios and enhancing the flexibility and applicability of VFL data reconstruction attacks.
@@ 3.1 System Model
where $\mathcal{D}$ is the training dataset, $\mathcal{L}$ is the loss function, and $\Theta=\{\theta_{1},\cdots,\theta_{K};\theta_{top}\}$ are VFL model parameters.
@@ 3.2 Threat Model
Adversary’s objective. The goal of the attack is to acquire the private data $x^{\text{target}}$ used by $\mathbb{P}_{\text{target}}$ in VFL training. This objective is quite ambitious, fundamentally challenging privacy preservation in VFL. Naturally, this is also very difficult to achieve, and nearly impossible in some practical scenarios. Therefore, the goal of the attack may be reduced to acquiring as much information about $x^{\text{target}}$ as possible, for example obtaining information about a specific column in $x^{\text{target}}$.
Adversary’s capacity. We assume that the attacker $\mathbb{P}_{\text{adv}}$ strictly follows the VFL protocol and cannot disrupt the normal operations on $\mathbb{P}_{\text{target}}$; therefore, $\mathbb{P}_{\text{adv}}$ has no access to the bottom model at $\mathbb{P}_{\text{target}}$. Since $\mathbb{P}_{\text{adv}}$ is on the active party, it can obtain the intermediate features $H^{1},\cdots,H^{K}$ from all participants including $\mathbb{P}_{\text{target}}$. In different VFL scenarios, $\mathbb{P}_{\text{adv}}$ may have the following capabilities: possesses an auxiliary dataset with an identical and independent distribution (i.i.d.) as $x^{\text{target}}$, make inference queries to the $\mathbb{P}_{\text{target}}$’s model, or obtain a minimal amount of private data samples used by $\mathbb{P}_{\text{target}}$ for local training.
@@
The adversary may launch a variety of attacks for data reconstruction based on the different capabilities that it has. Table 1 lists the attack scenarios and the required adversary capabilities. In the Query Attack scenario, the attacker possesses an i.i.d. auxiliary dataset and can make queries to $\mathbb{P}_{\text{target}}$. The Data Passive Attack only requires the attacker to own i.i.d. dataset as the target’s private data, while the Isolated Query Attack assumes the attacker can make queries to $\mathbb{P}_{\text{target}}$ but has no auxiliary dataset. In the Stealth Attack scenario, the attacker can neither make queries nor has i.i.d. data but has obtained a minimal amount of the target’s private data used in training its bottom model.

### v2

# UIFV: Data Reconstruction Attack in Vertical Federated Learning

@@ Abstract
Vertical Federated Learning (VFL) enables collaborative machine learning without the need for participants to share their raw private data. However, recent studies have uncovered privacy risks, where adversaries might reconstruct sensitive features through data leakage during the learning process. Although existing data reconstruction methods are effective to some extent, they exhibit limitations in VFL scenarios, as initiating an attack requires meeting more stringent conditions. To gain a comprehensive understanding of the risks of data reconstruction in VFL, this paper proposes a unified framework, the Unified InverNet Framework in VFL (UIFV), for data reconstruction under realistic black-box threat models. Within the UIFV framework, we consider four attack scenarios, strictly adhering to VFL protocols to maintain confidentiality. Experiments on four datasets show that our methods significantly outperform state-of-the-art techniques in terms of applicability and attack precision. Our work reveals severe privacy vulnerabilities within VFL systems that pose real threats to practical VFL applications, thus confirming the necessity of further enhancing privacy protection in the VFL architecture. Overall, this paper provides a thorough analysis of the risks of data reconstruction in VFL and offers important guidance to enhance the security of VFL deployments.
@@
journal: Pattern Recognition
[…]
[1]organization=School of Computer Science, Fudan University, city=Shanghai, postcode=200433, country=China
[…]
[2]organization=Shanghai Blockchain Engineering Research Center, city=Shanghai, postcode=200433, country=China
[…]
[3]organization=Institute of Financial Technology, Fudan University, city=Shanghai, postcode=200433, country=China
[…]
[4]organization=Information Sciences & Technology, Pennsylvania State University, city=PA, postcode=16802, country=USA
[…]
[5]organization=School of Software, Nanjing University of Information Science and Technology, city=Nanjing, postcode=210044, country=China
@@ 1 Introduction
In today’s field of artificial intelligence (AI), the integration of federated learning is regarded as a truly transformative strategy. This unique machine learning approach distributes the model training process across multiple devices and subsequently combines the model updates from these individual devices into a comprehensive global model. This process achieves a delicate balance by protecting data privacy while utilizing large-scale datasets for effective model training and optimization. Federated learning circumvents the need for direct access to or transmission of raw data, thereby significantly reducing the risk of data breaches. At the same time, the model’s ability to train on large datasets greatly enhances its performance and generalization capabilities. Now, federated learning is increasingly applied to real-life applications such as mobile keyboard prediction[1], healthcare[2], and purchase recommendations[3].
Figure 1: Data partitioning of HFL and VFL. (a) Horizontal partitioned data. (b) Vertical partitioned data.
Within the framework of federated learning, based on the distribution characteristics of local data, it is mainly divided into two scenarios: Horizontal Federated Learning (HFL) and Vertical Federated Learning (VFL), as shown in Fig. 1. In HFL, the local datasets of data owners have almost no intersection in the sample space, but there is a significant overlap in the feature space. In contrast, in VFL, local datasets have a large intersection in the sample space, but little overlap in the feature space. Each of these federated learning types has its applicable scenarios and advantages. VFL is particularly suitable for situations where different institutions hold different feature data of the same set of users; for example, in the financial sector, one institution may have the credit history of users, while another may have their transaction data. Through VFL, these institutions can collaborate to build more accurate risk assessment models without the need to directly exchange sensitive data. VFL has found wide applications in fields such as finance and healthcare.
Despite the advantages of VFL in protecting private data, recent studies have shown that it may still face risks of data privacy breaches [4], especially through data reconstruction attacks. Such attacks reconstruct the original features of the training dataset by analyzing intermediate data during the VFL process, potentially leading to sensitive information leaks.
@@
Fig. 2 shows a VFL application scenario that is vulnerable to data reconstruction attacks [5]. In this scenario, a bank and a fintech company participate in a VFL for credit analysis, with each entity possessing a subset of user attributes. An assessment agency uses the labels of users to coordinate the training of the VFL model. However, the assessment agency wants to acquire the fintech company’s private data and take it for its own use, so it launches a data reconstruction attack on the model owned by the fintech company. Without violating the VFL protocol, it uses the intermediate data from the VFL to reconstruct the private data owned by the bank, leading to the leakage of the fintech company’s customer information and posing a significant security risk to the actual application of VFL.
Existing data reconstruction attacks on VFL, such as the generative regression network (GRN) method [6] and the gradient-based inversion attack (GIA) method [7], largely draw from methods used in HFL (Horizontal Federated Learning), with a core focus on utilizing model information. However, these methods have certain limitations. For example, the GRN method is primarily based on a white-box scenario, where the attacker reconstructs input data by accessing the passive party’s model to calculate gradients in a multi-layer neural network. However, since passive parties in VFL typically do not share their models, this method is significantly constrained in practical applications. To address this limitation, the GIA method introduces a black-box attack scenario, where a proxy model is constructed, and data reconstruction is achieved by optimizing on this shadow model. Although the GIA method alleviates some limitations of the white-box scenario, it imposes strict requirements on model types, such as being applicable only to logistic regression (LR) models or neural networks without nonlinear activation functions in the output layer. Additionally, the GIA method requires individual optimization for each reconstruction, which further restricts its applicability and practicality due to its operational complexity and efficiency bottlenecks.
It is crucial to go beyond the existing attack methods and explore a broader range of attack scenarios in order to gain a more thorough understanding of the potential threats to data privacy in VFL. The goal of this paper is to fully consider real-world attack scenarios in practical VFL environments and develop effective attack methods for different scenarios. To accomplish this, we have completely abandoned the traditional approach that relies on gradient information or model information and instead have opted to directly utilize the intermediate feature data generated in the VFL framework for attacks. This method constructs an inverse net (InverNet) to effectively extract original data information from the intermediate features output from the target’s model. Following this strategy, We have developed an innovative attack framework, called Unified InverNet Framework in VFL (UIFV), that is applicable to a variety of VFL scenarios with different adversary capabilities. UIFV overcomes the dependency on gradient information or model information of traditional attacks in VFL, thereby providing a more flexible and effective means for data reconstruction attacks in complex VFL environments.
@@
  • 1.
    Exploration of VFL Data Reconstruction Risks: We thoroughly investigate the risks of data reconstruction in VFL, analyze potential threats in different attack scenarios, and provide important insights for future defense measures.
  • 2.
@@ 2.1 Vertical Federated Learning
The participants in a VFL framework consist of an active party and some passive parties [8]. Each passive party owns a set of data features that are fed into a model (called a bottom model) for local training. The active party holds the label information and a top model in addition to its own feature set and bottom model. The active party coordinates the training process by concatenating the bottom model outputs as the input to the top model. This structure enables collaborative model training without the need to share original data, thus preserving the privacy of sensitive data on the participants.
As an important branch of federated learning, participants in VFL typically include one active party and several passive parties. Each passive party has a set of data features and trains a local bottom model. The active party, in addition to having its own feature set and bottom model, also holds label information and a top model. The active party coordinates the entire training process by passing the output of the bottom models as input to the top model. This structure enables collaborative model training without sharing raw data, thereby protecting the sensitive data privacy of participants.
VFL has extensive application prospects in fields such as finance, advertising, and healthcare. Kang et al.[9] proposed a privacy-preserving VFL framework designed for financial applications, significantly improving the performance of credit loans. Li et al.[10] designed a label-protecting VFL framework that enhances advertising conversion rates. Fu et al.[11] made progress in the study of ductal carcinoma in situ (DCIS) using a VFL framework[12]. This paper provides a formal definition of VFL in section 3.1.
2.2 Attacks in VFL
Although VFL has made significant progress in practical fields such as finance and advertising, its potential security vulnerabilities have raised widespread concerns, especially threats from within the VFL system, where one or more participants may attempt to attack others. The internal security issues of VFL can be divided into two main categories: one is attacks that disrupt the normal operation of VFL. For example, Liu et al.[13] used Projected Gradient Descent[14](PGD) and feature flipping attacks to poison VFL predictions, and Chen et al.[15, 16] implanted backdoors during the VFL training phase, allowing passive parties to arbitrarily control prediction results during the inference phase. The other category involves attempts to obtain data from other participants, particularly private features or label information. Research on label attacks is relatively extensive. Li et al.[10] used direction and norm scoring methods to infer server labels based on distribution differences between positive and negative samples, and Zou et al.[17] stole server labels through gradient inversion. In comparison, data reconstruction attacks pose a more serious threat in VFL, as attackers can reconstruct the original input data of other parties by analyzing intermediate gradients or model parameters, severely threatening data privacy. Existing studies, such as the Generative Regression Network (GRN) method[6] and the Gradient-based Inversion Attack (GIA) method[7], have limited application in VFL scenarios, necessitating a more comprehensive analysis of the risks associated with VFL data reconstruction.
@@ 2.3 Data Reconstruction Attack
Method Attack Type Requires Requires Needs Model Modifies Requires Training Training/Inference Inference Gradient Structure and Model Non-IID data Query Query Parameters Architecture data Support Support DGL[18] Gradient-based - - - - Training SQR[19] - - - - Training CPA[20] - - - - Training LOKI[21] - - - Training GRN[6] Model - - - - - Training/Inference GIA[7] information-based - - $\circ$ - - Training/Inference Ginver[22] Feature-based - $\circ$ - - - Training/Inference UIFV $\circ$ - - - $\circ$ Training/Inference
Table 1: Comparison of Data Reconstruction Attack Methods and Their Requirements (✓indicates the presence of a specific requirement or feature, - indicates that the requirement is not needed, and $\circ$ denotes optional or partially required conditions. refers to whether the method supports only using a small amount of real leaked samples to assist the attack.)
2.3.1 Gradient-Based Methods
Gradient-based methods achieve data reconstruction by leveraging the gradients generated during the training process of machine learning models. These methods require the attacker to access both the model and the gradients of the target data. Currently, these methods can be categorized into two main types: The first type, represented by the pioneering work Deep Gradient Leakage (DGL) [18], focuses on initializing virtual data and calculating its gradients to achieve reconstruction. The attack process can be represented by the following equation:
$\hat{x}=\arg\min_{x}\|\nabla L(x,y,\theta)-\nabla W\|^{2}$ (1)
Here, $\nabla L(x,y,\theta)$ represents the generated gradients, $\nabla W$ is the true gradients, $x$ is the virtual data, $\hat{x}$ is the reconstructed data, $L$ is the loss function, $y$ is the label, and $\theta$ is the model parameter. Recent studies, such as [23] and [19], have further extended this approach with remarkable results. For instance, [23] proposed TabLeak, a method specifically designed for tabular data, while [19] used a tensor decomposition approach to reconstruct private samples through a single gradient query (In this paper, we refer to this method as SQR).
The second type, represented by the LOKI approach in [21], primarily targets attacks on linear models. These methods directly compute gradients to recover the original data. The attack process can be described by the following equation:
$\hat{x}=\frac{\delta L}{\delta W^{i}}/\frac{\delta L}{\delta B^{i}}$ (2)
Here, $\frac{\delta L}{\delta W^{i}}$ is the weight gradient, $\frac{\delta L}{\delta B^{i}}$ is the bias gradient of neuron $i$, and $x$ is the data that activates neuron $i$.
However, gradient-based methods face a significant limitation: their performance degrades rapidly as the batch size increases, due to the mixing of gradients from different samples. To address this issue, [20] employed Independent Component Analysis (ICA) to separate independent update signals and proposed the CPA method, successfully enabling reconstruction even with larger batch sizes. Additionally, [24] systematically described how attackers could exploit gradient information in federated learning and provided detailed steps for implementing these attacks. It is worth noting that these gradient-based methods are primarily applicable to HFL architectures. In VFL, where access to gradient information is limited, these methods face substantial challenges.
2.3.2 Model Information-Based Methods
Model-based reconstruction methods leverage the internal parameters of machine learning models to recover original training data. Unlike gradient-based methods, these approaches do not require access to gradient information during the training process but instead rely on access to the model and its internal parameters. The core idea is to initialize virtual data and compute its output on the model, optimizing it by comparing against the target output. The reconstruction process can be expressed as:
$\hat{x}=\arg\min_{x}\|f(x,\theta)-H\|^{2}$ (3)
where $H$ is the true output of the target data on the model $f$. [6] proposed the GRN method to recover the passive party’s data in VFL, which requires access to the passive party’s model and its parameters. Building on GRN, [7] introduced the GIA method, which uses a small amount of known auxiliary data and their confidence scores to construct a shadow model that simulates the passive party’s model, enabling data reconstruction without direct access.
Essentially, these methods rely on the internal parameters of the model and adopt a white-box approach to optimize virtual data. However, due to concerns over data privacy and security, participants in VFL frameworks are often reluctant to share model details, making it highly challenging to access the passive party’s model.
2.3.3 Feature-Based Methods
Feature-based methods leverage model outputs, such as Shapley values or intermediate features, to reconstruct data. These approaches do not require knowledge of model parameters or gradients. The core idea is to establish a relationship between the model’s output features and its input, aiming to reconstruct the input data based on the output features. The attack process can be described as:
$\hat{x}=g(H,\hat{\theta}_{g})\quad\text{where}\quad\hat{\theta}g=\arg\min{\theta_{g}}\left\|g(H,\theta_{g})-x\right\|^{2}$ (4)
Here, $g$ is a function that maps $H$ back to $x$, and $\theta_{g}$ represents the parameters of the function $g$. In traditional machine learning, [25] was the first to reveal the risks of feature inference attacks in model explanations based on Shapley values, demonstrating that current explanation methods can lead to privacy leakage.
In the field of split learning, some studies have also proposed attack methods targeting this issue. However, these methods have certain limitations. For example, the methods proposed by He et al. in [26, 27] heavily rely on auxiliary datasets to reconstruct user inputs in black-box settings. Additionally, the Ginver attack proposed in [22] relies on Inference Queries, requiring one query for each gradient update of $g$. It is noteworthy that these methods have not yet been validated in VFL architectures. Our method, UIFV, is specifically designed to address the characteristics of VFL. It innovatively treats the passive party model in VFL as a black box and further reduces the prerequisites for launching attacks through two modules: Data Preparation and Model Preparation. This design significantly broadens the application scenarios of VFL data reconstruction attacks while enhancing their flexibility and adaptability.
@@ 3.1 System Model
Figure 3: VFL Architecture Diagram with Three Participants
@@
where $\mathcal{D}$ is the training dataset, $\mathcal{L}$ is the loss function, and $\Theta=\{\theta_{1},\cdots,\theta_{K};\theta_{top}\}$ are VFL model parameters. A VFL framework with three participants (one active party and two passive parties) is illustrated in Fig. 3.
@@ 3.2 Threat Model
Adversary’s objective. The goal of the attack is to acquire the private data $x^{\text{target}}$ used by $\mathbb{P}_{\text{target}}$ in VFL training. This objective is quite ambitious, fundamentally challenging privacy preservation in VFL. Naturally, this is also very difficult to achieve and nearly impossible in some practical scenarios. Therefore, the goal of the attack may be reduced to acquiring as much information about $x^{\text{target}}$ as possible, for example, obtaining information about a specific column in $x^{\text{target}}$.
Adversary’s capacity. We assume that the attacker $\mathbb{P}_{\text{adv}}$ strictly follows the VFL protocol and cannot disrupt the normal operations on $\mathbb{P}_{\text{target}}$; therefore, $\mathbb{P}_{\text{adv}}$ has no access to the bottom model at $\mathbb{P}_{\text{target}}$. Since $\mathbb{P}_{\text{adv}}$ is on the active party, it can obtain the intermediate features $H^{1},\cdots,H^{K}$ from all participants, including $\mathbb{P}_{\text{target}}$. In different VFL scenarios, $\mathbb{P}_{\text{adv}}$ may have the following capabilities: possesses an auxiliary dataset with an identical and independent distribution (i.i.d.) as $x^{\text{target}}$, make inference queries to the $\mathbb{P}_{\text{target}}$’s model, or obtain a minimal amount of private data samples used by $\mathbb{P}_{\text{target}}$ for local training.
@@
The adversary may launch a variety of attacks for data reconstruction based on the different capabilities that it has. Table 2 lists the attack scenarios and the required adversary capabilities. In the Query Attack scenario, the attacker possesses an i.i.d. auxiliary dataset and can make queries to $\mathbb{P}_{\text{target}}$. The Data Passive Attack only requires the attacker to own an i.i.d. dataset as the target’s private data, while the Isolated Query Attack assumes the attacker can make queries to $\mathbb{P}_{\text{target}}$ but has no auxiliary dataset. In the Stealth Attack scenario, the attacker can neither make queries nor has i.i.d. data but has obtained a minimal amount of the target’s private data used in training its bottom model.